# Confidential Federated Computations

### v1

@@ 1 Introduction
In this paper, we introduce a new system architecture for federated learning and analytics [52], designed to enable confidentiality of server-side computations and provide externally verifiable privacy properties. In short, a long lived, TEE hosted service holds the keys to uploaded encrypted data from client devices, and controls access to that data by TEE-hosted data processing pipelines via data access tracking and policy enforcement. TEE hosted processing is subject to confidentiality, integrity, and external verifiability. The system enables a fleet of participating devices to keep control of their data, with individual uploads to the system tagged with a complete and closed set of computations in which the data may ever participate. Participating devices may further verify the privacy properties of the algorithms (e.g. a federated learning algorithm), as well as the properties of any unencrypted values released by the algorithm to the service provider (e.g., a specific DP guarantee). All uploaded messages and all intermediate results are encrypted with keys that are inaccessible to the service operator, so long as the TEEs’ confidentiality and integrity properties hold. Our system design admits future opportunities to integrate SMPC as a defense against TEE failure or leakage, where applicable.
@@ 2.1 Potential Adversaries and Capabilities
A malicious platform operator might instead simulate many end users, using known data for each end user. For example, if the analysis is intended to be a simple sum and the data access policy requires a certain number of users’ data to be aggregated in that sum, then a malicious platform operator might include a single real users’ data while simulating the rest of users as all holding the value zero—thereby producing a sum that is equal to the targeted user’s data. These Sybil attacks [29] can make it particularly difficult to provide verifiable proof of privacy in the face of potentially malicious platform operators. We require a system that can protect against Sybil attacks, for example by ensuring that an end user submitting their data to the system can verify that the analysis will be differentially private even if all other end users are actually fake or controlled by an adversary.
@@ 3.2 Trusted Computing Specific Design Principles
    Limits the attack surface of TEE hosted software.
@@
    Proving what source code is being run in a TEE, requiring remote attestation to client devices and their users on what binaries are being run in a TEE, with the hardware (e.g. CPU) manufacturer being the root-of-trust. We provide more detail on our approach in Section 3.6.
@@ 3.3 Overview
Separately, a data processing pipeline such as a Map-Reduce style histogram computation is executed on the server. This pipeline orchestrates (spins up, chains) TEE hosted data transformations to operate on the uploaded private data; these workloads request decryption keys from the ledger which may approve or deny the request subject to the access policy cryptographically bound to the data, and records the usage. Communication between these workloads is also encrypted. Finally, a data transformation workload may release aggregate results to untrusted space subject to differential privacy guarantees.
@@ 3.4 Trusted Infrastructure Primer
Oak is implemented as a split architecture, in which only the subset of the logic that needs to be trusted runs in the TEE (and therefore contributes to the TCB) and has access to the decryption keys, while the rest of the logic that only deals with encrypted data and does not have access to the decryption keys runs outside of the TEE. The trusted and untrusted components communicate with each other over a communication channel in the form of a UDS socket. This allows better separation of concerns, and only granting privileges to the minimal amount of logic that requires it.
@@
  • endorsements: additional data used to establish the trustworthiness of the evidence (e.g., intermediate certificates, signatures, inclusion proofs); produced or cached by the untrusted host running the TEE.
  • reference values: the policy that determines the acceptable range of values that the client trusts (e.g., TCB version number, individual measurements, etc.); usually hardcoded in the client code, or obtained over some trustworthy channel.
@@ 3.5 Ledger
The ledger is the TEE-hosted application that enables the device, after verifying the identity of software in the TEE, to reason about the full set of downstream uses in which any uploaded data may participate. Conceptually, the ledger combines a key store (holding keys needed to decrypt any uploaded data) with a stateful policy engine (that enforces that keys are given out only to other TEE-hosted applications with approved access).
In more technical terms, the ledger is a key unwrapping service that enforces stateful use-based privacy policies. This service allows secure asynchronous communication between data producers and consumers using pre-existing channels, such as uploads with HTTP POST or processing with Apache Beam [59]. At a high level, a publisher—in our case, an end user device or subsequently a TEE in a data processing pipeline—encrypts its data using authenticated encryption with associated data (AEAD) with an ephemeral key, wraps the ephemeral key with a ledger-generated public key using single-shot hybrid public key encryption (HPKE), then transmits the encrypted data and wrapped key using a channel of its choice. To decrypt the data, a consumer—in our case, a TEE in a data processing pipeline—provides the wrapped key and the identity of its TEE hosted data transformation; if the consumer is authorized under the data’s access policy, the ledger unwraps the key and returns it to the TEE hosted data transformation over a secure channel. Unless that transformation is a terminal step in the data processing pipeline, it will then act as a data producer and re-encrypt the processed data before exposing it outside the TEE (see Section 3.7). The ledger maintains multiple HPKE keys so that keys can be rotated without turning up or down servers; these HPKE keys are signed by an application-level key in the Oak evidence to bind them to the TEE instance that generated them.
@@ 3.5.1 Data Access Policies
Data access policies specify the sequence of TEE hosted transformations that can access the data; the access policy is indelibly bound to the corresponding encrypted data by including it in the associated data (AD) used when wrapping the encrypted data’s encryption key with HPKE. The access policy can be represented as a directed graph where each edge describes access conditions such as the specific TEE hosted binary, the other components in the TEE (e.g. kernel and system image), and any privacy- or security-relevant aspects of the binary’s configuration (e.g. $(\varepsilon,\delta)$ values for a DP aggregator). By combining the specific set of binaries that are able to access the data with knowledge of their source code, it is possible to reason about the overall server-side usage of the data.
Figure 2: An example access policy for the pipeline in figure 1 allowing the initial data to be processed first by a TEE that performs a SELECT operation, then by a TEE that implements a DP sum with $\varepsilon\leq 1$ and $\delta\leq 10^{-15}$. Each stage may only access the data once, and no other access is allowed.
@@ 3.5.2 Limiting Data Retention via Cryptographic Erasure
HPKE keys generated by the ledger have expiration times after which the ledger will forget about the HPKE key and all associated access records—thus crypto-erasing all uploaded data protected by that key. However, as we will describe in Section 5.5, accessing a secure time source from within a TEE is difficult, especially at high QPS. So instead of using a secure time source, the ledger uses an insecure but monotonically increasing time source that is exposed to clients in the issued_at and expiration times of all generated HPKE keys. Client devices will not upload data if the ledger’s time is too skewed: if the ledger’s clock is running significantly behind the current time, its HPKE keys will already be expired, and if the ledger’s clock is running significantly ahead of the current time, the issued_at time will not yet be reached. See Section 5.5 for potential future improvements in this space.
@@ 3.6.2 Ledger Attestation Verification and Uploading Encrypted Results
    Verifying attestation evidence, ensuring that it is rooted in a hardware root of trust such as AMD SEV-SNP, and that the stage0 firmware, kernel, and application binary hashes are all valid and that the provided encryption key was produced by the attested application binary. A valid binary hash is one that occurs in an out-of-band provided allowlist of known hashes, or one that has been endorsed by both Google and a 3rd party such as Rekor [2]. For the time being, we are using allowlist-based binary verification, but we expect to move to an endorsement-based approach in the future.
@@
    Validating the data access policy that was provided alongside the task assignment. Currently we allow only a predetermined set of access policies, but we expect to support more advanced validation processes in the future.
If both checks pass the client prints an attestation verification record containing the attestation evidence and data access policy (see next section). On Android, this record is currently sent to logcat [1].
The client then encrypts its data using the encryption key and the HPKE-based process described earlier, and uploads the encrypted data to a server-specified URI, at which point the encrypted data can be further processed server-side with TEE-hosted processing, subject to the ledger limiting access according to the data access policy as described earlier.
@@ 3.6.3 Enabling External Verifiability of the Data Processing
By printing out the attestation verification record, we make it possible for someone (e.g. a security researcher) to gather such logs from their own device, and to redo the same verification steps that the client performed. Inspecting the logged attestation evidence allows one to verify exactly which ledger binary generated the public key, and therefore which binary is protecting access to the corresponding decryption key. The evidence also proves that the ledger is running in a TEE protected by a hardware root-of-trust like AMD SEV-SNP. By inspecting the data access policy one can determine exactly how the uploaded data will be processed. See Section 3.8 for more details.
Note that we expect this approach to evolve in the future, e.g. by logging the attestation verification record to more suitable and accessible locations than the logcat buffer, and by making it easier to trigger one’s device to participate in the currently available confidential computations.
@@ 3.7 Data Processing
Theoretically, the entirety of the data might be processed in a single TEE; however, for the type of scale of workloads we aim to support, this would be infeasible in terms of compute and memory, and suggests the use of a graph of TEEs for parallel and pipelined processing of data. Our architecture distinguishes between orchestration, which only ever operates on encrypted data and thus can run on hardware without confidential computing protections enabled, and data transformations, which require decrypting the data and thus must run in TEEs. We define a data processing workload as consisting of untrusted infrastructure orchestrating a graph of data transformations that run in TEEs.
@@ 3.7.3 Proprietary Functions and User-Level DP
Note that this observation only holds if the untrusted function has limited power: the function must only execute on one user’s data at a time and produce a single output for that user. Consider an untrusted function that rather than taking an input from one user and producing one output, takes inputs from N users and produces N outputs. An adversarial version of this function could produce N outputs that all correspond to one of the input users, breaking the DP algorithm’s assumption that each input corresponds to a different user. The same attack could be carried out if the untrusted function has the ability to save state between executions on inputs from different users. Thus, when writing trusted applications that execute proprietary data transformations, the externally verifiable open-source code must ensure that state is cleared between executions of the proprietary code.
@@ 3.8 OSS Repositories
The Confidential Federated Compute [38] repository contains the source code for the ledger binary and a number of data processing pipeline binaries, more detailed documentation of their implementations, and instructions for mapping an attestation verification record logged by the client library to the various binaries that record attests to (e.g. to map an application-layer binary hash in the attestation evidence to a specific, reproducibly buildable binary).
@@ 4.1 Differentially Private Heavy Hitters (PHH)
Note that the server-side workload definition specifies a privacy budget in terms of epsilon and delta ($\varepsilon$, $\delta$), which is the maximum privacy loss that the server-side workload will incur. The platform will match the query against the required data access policies associated with the user data being accessed (not shown). Before releasing any uploaded client data to the server-side processing pipeline, the platform confirms that the server-side workload definition conforms with the access policy. In this example, a data access policy requiring $(\varepsilon^{*},\delta^{*})$-DP or stronger would permit the query from Listing 1 to run as long as $1\leq\varepsilon^{*}$ and $10^{-8}\leq\delta^{*}$. The access policy can also limit how often the data is used, for example specifying that any uploaded client data will be processed at most once by a TEE hosted binary containing PHH Aggregation Cores.

### v2

@@ Abstract
Change history: - Version 2025-03: New Section 3.8 on enhanced external verifiability, and Section 4.3 on server-side federated learning. Various other clarifications. - Version 2024-04: First release.
@@ 1 Introduction
In this paper, we introduce a new system architecture for federated learning and analytics [56], designed to enable confidentiality of server-side computations and provide externally verifiable privacy properties. In short, a long lived, TEE-hosted service holds the keys to uploaded encrypted data from client devices, and controls access to that data by TEE-hosted data processing pipelines via data access tracking and policy enforcement. TEE-hosted processing is subject to confidentiality, integrity, and external verifiability. The system enables a fleet of participating devices to keep control of their data, with individual uploads to the system tagged with a complete and closed set of computations in which the data may ever participate. Participating devices may further verify the privacy properties of the algorithms (e.g. a federated learning algorithm), as well as the properties of any unencrypted values released by the algorithm to the service provider (e.g., a specific DP guarantee). All uploaded messages and all intermediate results are encrypted with keys that are inaccessible to the service operator, so long as the TEEs’ confidentiality and integrity properties hold. Our system design admits future opportunities to integrate SMPC as a defense against TEE failure or leakage, where applicable.
@@ 2.1 Potential Adversaries and Capabilities
A malicious platform operator might instead simulate many end users, using known data for each end user. For example, if the analysis is intended to be a simple sum and the data access policy requires a certain number of users’ data to be aggregated in that sum, then a malicious platform operator might include a single real users’ data while simulating the rest of users as all holding the value zero—thereby producing a sum that is equal to the targeted user’s data. These Sybil attacks [31] can make it particularly difficult to provide verifiable proof of privacy in the face of potentially malicious platform operators. Properly implemented, central differential privacy protects each summand by clipping updates and adding noise, but this guarantee only protects against Sibyl attacks when an end user submitting their data to the system can verify that the analysis will be differentially private even if all other end users are actually fake or controlled by an adversary. Our system is designed to provide this property.
@@ 3.2 Trusted Computing Specific Design Principles
    Limits the attack surface of TEE-hosted software.
@@
    Proving what source code is being run in a TEE, requiring remote attestation to client devices and their users on what binaries are being run in a TEE, with the hardware (e.g. CPU) manufacturer being the root-of-trust. We present two approaches in Section 3.6.3 and Section 3.8.
@@ 3.3 Overview
Separately, a data processing pipeline such as a Map-Reduce style histogram computation is executed on the server. This pipeline orchestrates (spins up, chains) TEE-hosted data transformations to operate on the uploaded private data; these workloads request decryption keys from the ledger which may approve or deny the request subject to the access policy cryptographically bound to the data, and records the usage. Communication between these workloads is also encrypted. Finally, a data transformation workload may release aggregate results to untrusted space subject to differential privacy guarantees.
@@ 3.4 Trusted Infrastructure Primer
Oak is implemented as a split architecture, in which only the subset of the logic that needs to be trusted runs in the TEE (and therefore contributes to the TCB) and has access to the decryption keys, while the rest of the logic that only deals with encrypted data and does not have access to the decryption keys runs outside of the TEE. The trusted and untrusted components communicate with each other over a communication channel in the form of a virtual network socket. This allows better separation of concerns, and only granting privileges to the minimal amount of logic that requires it.
@@
  • endorsements: additional data used to establish the trustworthiness of the evidence (e.g., intermediate certificates, signatures of binary hashes, transparency log inclusion proofs); produced or cached by the untrusted host running the TEE.
  • reference values: the policy that determines the acceptable range of values that the client trusts (e.g., TCB version number, individual binary hashes, signing keys with which binary hashes can be endorsed, etc.); usually hardcoded in the client code, or obtained over some trustworthy channel.
@@ 3.5 Ledger
The ledger is the TEE-hosted application that enables the device, after verifying the identity of software in the TEE, to reason about the full set of downstream uses in which any uploaded data may participate. Conceptually, the ledger combines a key store (holding keys needed to decrypt uploaded data and intermediate, unreleased state) with a stateful policy engine (that enforces that keys are given out only to other TEE-hosted applications with approved access).
Specifically, the ledger is a key unwrapping service that enforces stateful use-based privacy policies. This service allows secure asynchronous communication between data producers and consumers using pre-existing channels, such as uploads with HTTP POST or processing with Apache Beam [1]. At a high level, a publisher—in our case, a user device or subsequently a TEE in a data processing pipeline—encrypts its data using authenticated encryption with associated data (AEAD) with an ephemeral key, wraps the ephemeral key with a ledger-generated public key using single-shot hybrid public key encryption (HPKE), then transmits the encrypted data and wrapped key using a channel of its choice. To decrypt the data, a consumer—in our case, a TEE in a data processing pipeline—provides the wrapped key and the identity of its TEE-hosted data transformation; if the consumer is authorized under the data’s access policy, the ledger unwraps the key and returns it to the TEE-hosted data transformation over a secure channel. Unless that transformation is a terminal step in the data processing pipeline, it will then act as a data producer and re-encrypt the processed data before exposing it outside the TEE (see Section 3.7). The ledger maintains multiple HPKE keys so that keys can be rotated without turning up or down servers; these HPKE keys are signed by an application-level key in the Oak evidence to bind them to the TEE instance that generated them.
@@ 3.5.1 Data Access Policies
A data access policy specifies a specific sequence of TEE-hosted transformations that can access uploaded data. Devices indelibly bind their encrypted uploads to a data access policy that fully determines the closed set of processing steps the data may partake in, as enforced by the ledger. The data access policy is uploaded as associated data (AD) used when wrapping the encrypted data’s encryption key with HPKE. The access policy can be represented as a directed graph where each edge describes access conditions such as the specific TEE-hosted binary (and other components in the TEE including kernel and system image) as well as any privacy- or security-relevant aspects of the binary’s configuration (e.g. $(\varepsilon,\delta)$ values for a DP aggregator). By combining the specific set of binaries that are able to access the data with knowledge of their (open) source code, it is possible to reason about the overall server-side usage of the data.
Figure 2: An example access policy for the pipeline in figure 1 allowing the initial data to be processed first by a TEE that performs a SELECT operation, then by a TEE that implements a DP sum with $\varepsilon\leq 1$ and $\delta\leq 10^{-15}$. Each stage may only access the data once, and no other access is allowed. Data (gray) and transformations (green) with a “lock” icon are inaccessible to the system operator, assuming correct ledger and TEE functionality.
@@ 3.5.2 Limiting Data Retention via Cryptographic Erasure
HPKE keys generated by the ledger have expiration times after which the ledger will forget about the private key and all associated access records—thus crypto-erasing all uploaded data protected by that key. However, as we will describe in Section 5.5, accessing a secure time source from within a TEE is difficult, especially at high QPS. So instead of using a secure time source, the ledger uses an insecure but monotonically increasing time source that is exposed to clients in the issued_at and expiration times of all generated HPKE keys. Client devices will not upload data if the ledger’s time is too skewed: if the ledger’s clock is running significantly behind the current time, its HPKE keys will already be expired, and if the ledger’s clock is running significantly ahead of the current time, the issued_at time will not yet be reached. See Section 5.5 for potential future improvements in this space.
@@ 3.6.2 Ledger Attestation Verification and Uploading Encrypted Results
    Verifying ledger attestation evidence, ensuring that it is rooted in a hardware root of trust such as AMD SEV-SNP, and that the stage0 firmware, kernel, and application binary hashes are all valid and that the provided encryption key was produced by the attested application binary.
@@
    Validating the data access policy that was provided alongside the task assignment.
Upon successful completion of these checks, the client encrypts its data using the encryption key and the HPKE-based process described earlier, and uploads the encrypted data to a server-specified URI, at which point the encrypted data can be further processed server-side with TEE-hosted processing, subject to the ledger limiting access according to the data access policy as described earlier.
@@ 3.6.3 Enabling External Verifiability of the Data Processing
There are multiple ways in which a client can validate the ledger’s attestation evidence and the provided access policy. A simple approach uses an allowlist of hashes that is provided to the client device out-of-band. If the received attestation evidence and acccess policy are both allowlisted, then the client locally logs (in today’s system for Android, to logcat [3]) an attestation verification record containing the attestation evidence and data access policy.
By examining the attestation verification record from a device’s logs, a user would be able to manually follow the same steps as the device did to verify the identity of the ledger binary that generated the public key and the corresponding ledger implementation that protects access to the corresponding decryption key. The manually verified evidence can also prove that the ledger is running in a TEE protected by a hardware root-of-trust like AMD SEV-SNP. And by inspecting the data access policy, the full set of data processing steps and their corresponding implementations can be determined.
This simple approach requires gathering attestation verification records from devices one at a time. Its limitation is the infeasibility of instrumenting all devices that could possibly contribute data, which might (in the worst case) be necessary to discover the full set of access policies and binary endorsements served to and accepted by devices. In Section 3.8 we describe a more scalable approach to validating ledger attestation evidence and access policies based on a public transparency log.
@@ 3.7 Data Processing
Theoretically, the entirety of the data might be processed in a single TEE; however, for the scale of workloads we aim to support, this would be infeasible in terms of compute and memory, and suggests the use of a graph of TEEs for parallel and pipelined processing of data. Our architecture distinguishes between orchestration, which only ever operates on encrypted data and thus can run on hardware without confidential computing protections enabled, and data transformations, which require decrypting the data and thus must run in TEEs. We define a data processing workload as consisting of untrusted infrastructure orchestrating a graph of data transformations that run in TEEs. For many applications (as we will see in the next section), this provides a sufficient degree of verifiability while also allowing flexibility in data processing patterns and infrastructure. [^6] For example, central DP guarantees often require global consistency with respect to the ordering or grouping of incoming data. A user-level DP algorithm might, for instance, require that all of one user’s data be processed on a single node for joint clipping. Or a DP learning algorithm might require that gradients in later stages are applied in an order pre-determined by an early stage. In these cases, early data processing steps must communicate global consistency requirements to later steps, which can assert that the untrusted orchestrator provides data in the expected way. Thus, a malicious orchestrator would be able to slow or halt progress (a processing step’s assertions on data consistency fail) but would not be able to compromise the data processing pipelines’ anonymization guarantees beyond what they might learn about ordering.
@@ 3.7.3 Proprietary Functions and User-Level DP
This observation only holds if the untrusted function has limited power: the function must only execute on one user’s data at a time (optionally with other already-DP-protected inputs derived from other user’s data, like the current model iterate) and produce a single output for that user. Consider an untrusted function that rather than taking an input from one user and producing one output, takes inputs from N users and produces N outputs. An adversarial version of this function could produce N outputs that all correspond to one of the input users, breaking the DP algorithm’s assumption that each input corresponds to a different user. The same attack could be carried out if the untrusted function has the ability to save state between executions on inputs from different users. Thus, when writing trusted applications that execute proprietary data transformations, the externally verifiable open-source code must ensure that state is cleared between executions of the proprietary code.
@@ 3.7.4 Aggregation Cores
3.8 Enhanced External Verifiability
In Section 3.6.3 we described a simple allowlist-based mechanism through which client devices can validate and log ledger attestation evidence as well as the data access policies they are served, enabling a moderate level of verifiability for anyone instrumenting the device. We also established an important downside to this approach: that it doesn’t allow one to determine the full set of ledger binaries or access policies that any given device, not just a specific instrumented device, might accept.
This section describes a complementary mechanism which achieves an enhanced level of external transparency, by enabling anyone to determine the complete set of data processing steps that could apply to the data uploaded by any device, without having to instrument any devices at all. The approach augments the architecture described in the previous sections with cryptographic endorsements of binaries and access policies, which are then published to a transparency log.
Ledger and access policy endorsements
TEE binary endorsements—cryptographic signatures of the binary’s digest—are used to indicate to a client device that a binary has been signed by the service provider to indicate its correct operation. By endorsing binaries for each layer of the TEE-hosted ledger (firmware, kernel, application, etc.), devices can verify that the service provider has claimed that uploads’ decryption keys will be released only in accordance with those uploads’ access policies. And of course, this claim can be verified by anyone who examines (and reproducibly builds) the ledger binaries resulting in the same endorsed binary digest.
Access policy endorsements—cryptographic signatures over an access policy’s digest—are used to indicate to a client device that the policy has been signed by the service provider as appropriately privacy preserving. For example, an access policy endorsement could be used to indicate that only differentially private aggregates can be released to the service provider, a claim that can, in turn, be verified by examining (and reproducibly building) the binaries as associated with the processing steps referenced by the access policy.
Note that this architecture allows client devices to require ledger binaries or access policies to be endorsed by multiple signing parties. A client application that requires multiple endorsements could, for example, require that the ledger binary or processing steps be vetted by an independent auditor (not the platform software authors). While today’s production system does not currently use such configurations, we see this as promising area of future work.
Transparency logs
Transparency logs are tamper-evident structures to which data is published in a non-revertible, append-only manner. Violations of this append-only property are publicly detectable. Transparency logs can also generate for each logged entry an inclusion proof, or a cryptographic signature made by the transparency log operator that the entry has been included in the public log. Inclusion proofs allow devices to verify signed items are transparent without interacting with the transparency service, as long as one trusts the log operator’s handling of the inclusion proof signing key.
To enable public inspection of our ledger binaries and access policies, we publish the ledger binary and access policy endorsements described above to a transparency log. We use the Rekor [5] transparency log service for this purpose, part of the SigStore project and operated by an independent non-profit organization. The transparency log operator can thus be considered an adversary that is separate and non-overlapping with the other adversaries in our threat model (see Section 2).
Before accepting ledger attestation evidence and access policies through the endorsement verification process, client devices can also validate that endorsements also have a corresponding transparency log inclusion proof. In this way, external parties can monitor the transparency log service to determine the full of set of ledger binaries and data access policies that a client device could ever accept. Doing so requires only monitoring the log for new entries created by an endorsing key of interest.
Thus, external parties can determine all possible code paths through which uploaded client data might be processed. By inspecting the provenance (open-source code and build instructions) for all relevant TEE-hosted binaries (ledger and data processing steps) and access policies, anyone can follow along and validate or falsify the privacy properties of data processed in the system.
@@ 3.9 OSS Repositories
The Confidential Federated Compute [40] repository contains the source code for the ledger binary and a number of data processing pipeline binaries, documentation of their implementations, and instructions for mapping binary digests (logged in an attestation verification record produced by a client, or endorsed in transparency log entries) to the corresponding reproducibly buildable binaries and their source code.
@@
Scalable multi-TEE services
To prevent state from being lost when the ledger TEE goes down, its implementation in the Confidential Federated Compute repo leverages the utilities for multi-TEE distributed services implemented in the Trusted Computations Platform [42] repository.
@@ 4.1 Differentially Private Heavy Hitters (PHH)
Note that the server-side workload definition specifies a privacy budget in terms of epsilon and delta ($\varepsilon$, $\delta$), which is the maximum privacy loss that the server-side workload will incur. The platform will match the query against the required data access policies associated with the user data being accessed (not shown). Before releasing any uploaded client data to the server-side processing pipeline, the platform confirms that the server-side workload definition conforms with the access policy. In this example, a data access policy requiring $(\varepsilon^{*},\delta^{*})$-DP or stronger would permit the query from Listing 1 to run as long as $1\leq\varepsilon^{*}$ and $10^{-8}\leq\delta^{*}$. The access policy can also limit how often the data is used, for example specifying that any uploaded client data will be processed at most once by a TEE-hosted binary containing PHH Aggregation Cores.